OpenStudy (pulsified333):

5x-10y=25 24x+ay=120 What is this exceptional value for a?

OpenStudy (anonymous):

what the monkey does "exceptional value" mean?

OpenStudy (pulsified333):

no clue

Directrix (directrix):

The one value of "a" for which the system of equations does NOT have a unique solution. Exceptional Value as defined elsewhere.

OpenStudy (pulsified333):

so how would i solve for a?

Directrix (directrix):

I have a candidate for an answer.

Directrix (directrix):

I took this: 5x-10y=25 and divided each member by the common factor of 5. x - 2y = 5 which means x = 5 + 2y In the second equation, 24x+ay=120 I replaced x by (5 + 2y) as in solving the system by substitution.

Directrix (directrix):

which led to: 24(5 + 2y)+ay=120 120 + 48 y + ay = 120

Directrix (directrix):

The 120 terms subtracted off, leaving 48y + ay = 0

Directrix (directrix):

Factoring out y: y * (48 + a) = 0 Then the Zero Product Property y = 0 or 48 + a = 0 y = 0 or a = -48

OpenStudy (pulsified333):

you sir are a genius

Directrix (directrix):

Then, I put the -48 in for a in the original equation. 5x-10y=25 24x - 48y=120 and graphed the two equations. They were the same equation in disguise and the solution is a single line with infinitely many solutions. So, a = -48 is the exceptional value. http://www.wolframalpha.com/input/?i=5x-10y%3D25%2C++24x+-48y%3D120